Please assist me in solving the following system by substitution:
y = 2x - 8
4x + 3y = 1
My initial efforts seem to be missing at least one step. Thank you.

Answer by kvlsravya(8)

Step 1:
Let us assume the given system of equations as
y = 2x - 8 ---- (1)
4x + 3y = 1 ---- (2)
Step 2:
Substitute (1) in (2), we get
4x + 3(2x - 8) = 1
4x + 6x - 24 = 1
10x - 24 = 1
10x = 25
Dividing by 10 on both sides, we get
x = 5/2
Step 3:
Plugging in the obtained value of x in any of the equatons (1), (2), we get
y = -3
Step 4:
Hence,the solution to the given system of equations is x = 5/2 and y = -3.